OpenStudy (anonymous):

Twelve times the sum of eight and a number is -29. What is the number? On scratch paper, translate the statement above into an equation that you can solve to answer this question. Do not solve it yet. Use x as your variable. Then solve your equation for x.

OpenStudy (anonymous):

look. The statement "Twelve times the sum of eight and a number" is translated into equation this way: 12(8+x), where x is a number 12(8+x)=-29

OpenStudy (anonymous):

divide both sides by 12: 12(8+x)/12=-29/12 8+x=-29/12 subtract 8 from both sides: 8+x-8=-29/12 - 8 x=\[-29/12-8\]

OpenStudy (anonymous):

I did it this way 12(8+x)=-29 12 times 8 96+12x=-29 -96 12x=-125 divide 12 x=-10.4

OpenStudy (yoloshroom):

Hey again monicaj do u still need help on this?

OpenStudy (anonymous):

yes I still don't understand, because when I put -10.4 in for x it doesn't give me -29

OpenStudy (yoloshroom):

Lemme take a look :)

OpenStudy (yoloshroom):

10.4 is the number i got aswell.. hmm @Compassionate

OpenStudy (anonymous):

so is 10.4 correct then